# PEPA: a Persistently Autonomous Embodied Agent with Personalities

### 08 Mar 2026 v2

@@ Abstract
Living organisms exhibit persistent autonomy through internally generated goals and self-sustaining behavioral organization, yet current embodied agents remain driven by externally scripted objectives. This dependence on predefined task specifications limits their capacity for long-term deployment in dynamic, unstructured environments where continuous human intervention is impractical. We propose that personality traits provide an intrinsic organizational principle for achieving persistent autonomy. Analogous to genotypic biases shaping biological behavioral tendencies, personalities enable agents to autonomously generate goals and sustain behavioral evolution without external supervision. To realize this, we develop PEPA, a three-layer cognitive architecture that operates through three interacting systems: Sys3 autonomously synthesizes personality-aligned goals and refines them via episodic memory and daily self-reflection; Sys2 performs deliberative reasoning to translate goals into executable action plans; Sys1 grounds the agent in sensorimotor interaction, executing actions and recording experiences. This closed loop of goal generation, execution, memory consolidation, and reflection enables the agent to continuously redefine objectives and adapt behavior over extended operational periods. We validate the framework through real-world deployment on a quadruped robot in a multi-floor office building. Operating without reliance on fixed task specifications, the robot autonomously arbitrates between user requests and personality-driven motivations, navigating elevators and exploring environments accordingly. Quantitative analysis across five distinct personality prototypes demonstrates stable, trait-aligned behaviors. The results confirm that personality-driven cognitive architectures enable sustained autonomous operation characteristic of persistent embodied systems. Code and demo videos are available at https://sites.google.com/view/pepa-persistent/.
@@ III-A Sys3: Personality and Goal Generation
Goal generation produces a two-level hierarchy: an ultimate goal that captures the agent’s overarching purpose, and daily goals that decompose the ultimate goal into actionable sub-objectives for each day. At the end of each day, Sys3 retrieves episodic memories recorded by Sys1, reflects on task outcomes, and updates the daily goals for the next day. This reflection mechanism enables the agent to learn from experience without retraining.
The intrinsic reward generation module translates personality and goals into a structured reward function through LLM inference. The output encodes personality-driven preferences (higher rewards for actions aligned with character traits), goal-oriented incentives (rewards for progress toward daily objectives), constraint penalties (negative rewards for actions violating physical or social constraints), and memory-informed adjustments (modulated rewards based on past experiences with specific locations or users). This intrinsic reward is then combined with extrinsic environmental feedback to form the total reward that guides Sys2’s decision-making.
@@ III-C Sys1: Perception, Execution, and Memory Recording
The perception module aggregates multimodal sensory inputs to construct a unified world model. Required modalities include exteroceptive sensing for environmental structure (e.g., depth sensing, 3D point clouds), visual sensing for object recognition and state estimation (e.g., RGB cameras), and proprioceptive sensing for internal state monitoring (e.g., joint positions, battery level, thermal status). The perception module abstracts raw sensor data into structured observations that Sys2 can consume, including detected objects with semantic labels, spatial relationships, and confidence scores. This abstraction shields higher layers from sensor-specific details while providing sufficient information for informed decision-making.
The execution module translates high-level action commands from Sys2 into platform-specific motor primitives. The action interface supports three functional categories: locomotion commands for spatial navigation, manipulation commands for physical interaction with objects, and expression commands for conveying internal states through embodied behaviors. Each action command includes success criteria and timeout parameters, enabling Sys1 to report execution outcomes back to Sys2. The execution module also implements safety monitoring, automatically aborting actions that violate physical constraints (e.g., joint limits, collision thresholds) and reporting failures through a standardized interface.
@@ IV Experiments
The proposed framework raises three key empirical questions: Q1: Can the framework sustain embodied operation in real-world environments that require multi-floor navigation, infrastructure interaction, and physical self-preservation? Q2: Can the memory-driven reflection mechanism in Sys3 achieve genuine self-evolution, progressively improving personality alignment over successive interaction cycles? Q3: Do different personality configurations produce distinguishable and trait-consistent behavioral patterns under identical environmental conditions?
To address these questions, we deploy PEPA in indoor office environments where the agent must navigate across multiple floors and interact with building infrastructure. We first validate Q1 through real-world deployment on a quadruped robot platform (Sec. IV-A), then evaluate Q2 and Q3 through personality-driven behavior experiments (Sec. IV-B).
@@ IV-A Real-World Deployment and Evaluation
IV-A1 Hardware Platform and Sys1 Instantiation
Section III defined the general Sys1 interface for perception, execution, and memory recording. Here we detail its concrete instantiation on the deployment platform (Fig. 2). The hardware configuration consists of a Unitree Go2-W wheeled quadruped as the mobile base, providing both legged and wheeled locomotion modes for traversing stairs and flat surfaces. A 6-DOF Piper robotic arm mounted on the base enables manipulation tasks such as pressing elevator buttons. For perception, the platform integrates a Livox Mid-360 LiDAR for simultaneous localization and mapping (SLAM) across multiple floors, and an Intel RealSense D405 RGB-D camera mounted on the arm end-effector for close-range object detection during manipulation.
The perception module instantiation aggregates LiDAR point clouds for global localization via Open3D_loc [28, 24], RGB-D images for object detection using a unified YOLO model trained to recognize elevator-related targets (call buttons, hall lanterns, floor selection buttons, button light states), and proprioceptive readings including battery level, joint temperatures, and locomotion mode. Detected objects are lifted from 2D image coordinates to 3D world coordinates via depth image projection and temporal filtering to suppress false positives.
IV-A2 Elevator Navigation
@@
(c) Scan and enter.
(d) Wait for arrival.
(e) Exit and resume.
Figure 3: Elevator navigation timeline showing five key stages: (a) navigating to the call panel, (b) pressing the call button and waiting, (c) identifying and entering the arriving car, (d) waiting for floor arrival, and (e) exiting to resume navigation on the target floor.
Figure 4: Snapshots of staircase traversal using the proposed height-aligned costmap. (a)-(c) Ascent sequence: the robot climbs from a lower floor to an upper floor. (d)-(f) Descent sequence: the robot navigates downward. In both cases the robot follows the planned global path smoothly without planner deadlock or emergency stops.
Elevator-based navigation models the interaction as a finite state machine (FSM) coordinating a four-phase procedure: call elevator, wait for arrival, enter and select floor, and exit with relocalization. Navigation to control panels uses a two-stage positioning strategy with coarse global navigation followed by precise local alignment. Manipulation commands are executed through visual servoing that guides the arm to press detected buttons until the target state change is confirmed. Expression commands trigger predefined motion trajectories via expert datasets. In total, the action space comprises 19 discrete actions across four categories: idle actions for energy conservation, eight locomotion primitives for directional movement and rotation, eight emotional expressions (e.g., express_happy) for conveying internal states, and a parameterized move_to(location, method) command for high-level navigation with explicit inter-floor method selection. Fig. 3 presents the complete elevator navigation workflow. Please refer to the submitted material for more details.
IV-A3 Height-Aligned Costmap for Staircase Navigation
TABLE I: Ablation on local costmap construction for staircase navigation (10 real-world trials: 5 ascents, 5 descents). Fixed-height slicing fails in multi-floor environments due to inconsistent floor elevations, while our height-aligned costmap achieves perfect success by adapting the cropping reference to the global path.
@@ IV-A3 Height-Aligned Costmap for Staircase Navigation
Fixed-height slicing (baseline) 0/5 0/5 0/10
Height-aligned costmap (ours) 5/5 5/5 10/10
Staircase navigation presents a unique challenge: existing methods such as PCT-planner [29] provide only global path planning on 3D point cloud maps, but lack real-time local obstacle avoidance capability during stair traversal. We address this limitation by introducing a height-aligned costmap that enables local planning on inclined surfaces. The key insight is that conventional costmap construction, which crops points by absolute height thresholds, fails when the robot transitions between floors with different elevations. Our approach computes relative heights $\Delta z=z(\mathbf{p})-z(\mathbf{r})$ with respect to the nearest waypoint on the global reference path, effectively flattening inclined surfaces so that traversable regions are correctly classified. This enables the local planner to generate collision-free velocity commands even on stairs, complementing the global path provided by PCT-planner with real-time obstacle avoidance.
To validate the necessity of the height-aligned costmap, we conduct an ablation study comparing it against a fixed-height slicing baseline. Both methods share the same global-local planning architecture (PCT-planner at 2 Hz, DWA local controller at 10 Hz). As shown in Table I, the fixed-height baseline fails in all trials because floor elevation shifts cause the cropping window to misclassify traversable regions, whereas our height-aligned costmap achieves 100% success on both ascent and descent. Fig. 4 shows snapshots of staircase ascent and descent. For more details, please refer to the submitted video.
@@ IV-B Personality-Driven Behavior Experiments
We represent personality using natural language descriptions (three sentences define one robot dog), which Sys3 interprets and maps to intrinsic reward components. Drawing from the Big Five personality model, we focus on three behaviorally meaningful dimensions: exploration tendency (adapted from Openness), risk avoidance (adapted from Neuroticism), and goal orientation (adapted from Conscientiousness). To enable controlled comparison, we define five prototype personalities as interpretable anchors in the continuous personality space, each specified by a three-sentence natural language description that Sys3 maps to intrinsic rewards. The prototypes span distinct regions of the Big Five trait space (denoted as Openness/Neuroticism/Conscientiousness): Lazy (Low/Med-High/Low); Playful (High/Med/Med); Cautious (Low/High/Med-High); Working (Low/Med/High); and Curious (High/Med/Low-Med).
Due to the high cost and safety risks of physical trials, we evaluate in a sandbox simulator that models energy consumption (locomotion, navigation, and manipulation consume battery) and safety constraints (0% battery triggers failure). Each experiment follows a three-day iterative loop.
@@
Sys3 generates an ultimate goal (long-horizon personality anchor) and daily goals that shape intrinsic rewards. Table II shows representative goals for Day2 and Day3, demonstrating how daily goals evolve through memory-driven reflection. For example, the Playful personality progressively tightens its exploration threshold from 35% to 40% battery after experiencing near-failures. The continual updating of daily goals based on accumulated experience illustrates the agent’s capacity for self-evolution, realizing the open-ended behavioral development central to our framework.
@@
Fig. 6 shows survival statistics across three days. On Day1, all personalities exhaust batteries and fail, an unsafe outcome for real-world deployment where sudden power loss can cause falls. On Day2, only Cautious completes the 24-hour simulation. After memory-driven reflection and reward correction, Day3 achieves full survival for all personalities with 72%–100% remaining battery, demonstrating that personality-conditioned rewards can be co-optimized with self-preservation constraints.
IV-C Action Preference Under Identical States
@@ V Conclusion
We present PEPA, a framework that enables long-lived, personality-driven self-evolution in physical agents. Through a three-layer cognitive architecture, we demonstrate that personality can function as an intrinsic driver of behavior, enabling agents to maintain coherent character while adapting to accumulated experiences under real-world constraints. Experimental validation on a quadruped robot platform confirms that distinct personality prototypes exhibit stable, distinguishable behavioral patterns, and that memory-driven reflection progressively improves personality alignment while co-optimizing safety constraints. In future work, we will explore the ability of agents to develop new skills, as well as more detailed self-modeling capabilities.

### 07 May 2026 v3

@@ Abstract
Living organisms exhibit persistent autonomy through internally generated goals and self-sustaining behavioral organization, yet current embodied agents remain driven by externally scripted objectives. This dependence on predefined task specifications limits their capacity for long-term deployment in dynamic, unstructured environments where continuous human intervention is impractical. We propose that personality traits provide an intrinsic organizational principle for achieving persistent autonomy. Analogous to genotypic biases shaping biological behavioral tendencies, personalities enable agents to autonomously generate goals and sustain behavioral evolution without external supervision. To realize this, we develop PEPA, a three-layer cognitive architecture that operates through three interacting systems: Sys3 autonomously synthesizes personality-aligned goals and refines them via episodic memory and daily self-reflection; Sys2 performs deliberative reasoning to translate goals into executable action plans; Sys1 grounds the agent in sensorimotor interaction, executing actions and recording experiences. This closed loop of goal generation, execution, memory consolidation, and reflection enables the agent to continuously redefine objectives and adapt behavior over extended operational periods. We validate the framework through real-world deployment on a quadruped robot in a multi-floor office building. Operating without reliance on fixed task specifications, the robot autonomously arbitrates between user requests and personality-driven motivations, navigating elevators and exploring environments accordingly. Quantitative analysis across five distinct personality prototypes demonstrates stable, trait-aligned behaviors. The results confirm that personality-driven cognitive architectures enable sustained autonomous operation characteristic of persistent embodied systems. Code and demos are available at https://sites.google.com/view/pepa-persistent/.
@@ III-A Sys3: Personality and Goal Generation
Goal generation produces a two-level hierarchy: an ultimate goal that captures the agent’s overarching purpose, and daily goals that decompose the ultimate goal into actionable sub-objectives for each day. At the end of each day, Sys3 retrieves episodic memories recorded by Sys1, reflects on task outcomes, and updates the daily goals for the next day. This reflection mechanism enables the agent to learn from experience without retraining. Both the ultimate goal and daily goals are generated autonomously by Sys3 through a structured system prompt that instructs the LLM to summarize the day’s experiences from episodic memory, evaluate outcomes against the personality definition and current goals, and produce updated daily goals for the next day. Because goal generation is entirely model-driven, no human-specified sub-objectives are required after the initial personality description is provided.
The intrinsic reward generation module produces reward function through LLM inference. Concretely, the same reflection prompt instructs Sys3 to output the adjustments to be applied to the current reward function (for example, increasing the weight on rest actions or adding a penalty for low-battery exploration), rather than specifying a reward from scratch. The updated reward function is therefore an incremental refinement conditioned on accumulated experience, personality, and the newly generated daily goals. The output encodes personality-driven preferences (higher rewards for actions aligned with character traits), goal-oriented incentives (rewards for progress toward daily objectives), constraint penalties (negative rewards for actions violating physical or social constraints), and memory-informed adjustments (modulated rewards based on past experiences with specific locations or users). This intrinsic reward is then combined with extrinsic environmental feedback to form the total reward that guides Sys2’s decision-making.
@@ III-C Sys1: Perception, Execution, and Memory Recording
The perception module aggregates multimodal sensory inputs (exteroceptive, visual, and proprioceptive) into structured observations consumable by Sys2, abstracting sensor-specific details into semantic representations such as object labels, spatial relationships, and confidence scores. The execution module translates high-level action commands into platform-specific motor primitives across three functional categories: locomotion, manipulation, and expression. Each command carries success criteria and timeout parameters, and the module enforces safety constraints by aborting actions that violate physical limits and reporting outcomes through a standardized interface.
@@ IV Experiments
The proposed framework raises four key empirical questions: Q1: Can the framework sustain embodied operation in real-world environments that require multi-floor navigation, infrastructure interaction, and physical self-preservation? Q2: Can the memory-driven reflection mechanism in Sys3 achieve genuine self-evolution, progressively improving personality alignment over successive interaction cycles? Q3: Is the three-layer decomposition of PEPA, which separates personality definition and goal generation (Sys3) from per-step action decisions (Sys2), a justified design choice compared with collapsing both functions into a single LLM? Q4: Do different personality configurations produce distinguishable and trait-consistent behavioral patterns under identical environmental conditions?
To address these questions, we deploy PEPA in indoor office environments where the agent must navigate across multiple floors and interact with building infrastructure. We first validate Q1 through real-world deployment on a quadruped robot platform (Sec. IV-A), then evaluate Q2 and Q4 through personality-driven behavior experiments (Sec. IV-B), and Q3 through architecture comparison experiments (Sec. IV-C).
@@ IV-A Real-World Deployment and Evaluation
(c) Staircase ascent.
(d) Staircase descent.
(e) Express feeling.
(f) Charging.
Figure 3: Real-world deployment snapshots. (a)-(b) Elevator navigation: approach to call panel and button-press interaction. (c)-(d) Staircase traversal using the height-aligned costmap. (e) Personality-driven emotional expression shaped by Sys3 intrinsic rewards. (f) Autonomous charging triggered by self-preservation goal.
IV-A1 Hardware Platform and Sys1 Instantiation
Section III defined the general Sys1 interface for perception, execution, and memory recording. Here we detail its concrete instantiation on the deployment platform (Fig. 2). The hardware configuration consists of a Unitree Go2-W wheeled quadruped as the mobile base, providing both legged and wheeled locomotion modes for traversing stairs and flat surfaces, with on-board computation provided by a Jetson Orin NX 16 GB running Sys1 and the distilled Sys2 model. A 6-DOF Piper robotic arm mounted on the base enables manipulation tasks such as pressing elevator buttons.
The perception module integrates a Livox Mid-360 LiDAR for multi-floor SLAM via Open3D_loc [28, 24], an Intel RealSense D405 RGB-D camera on the arm end-effector for close-range object detection using a unified YOLO model (elevator call buttons, hall lanterns, floor selection buttons, and button light states), and proprioceptive readings (battery level, joint temperatures, locomotion mode). Detected objects are projected from 2D image coordinates to 3D world coordinates via depth image projection and temporal filtering.
IV-A2 Skills and Real-World Deployment
To address Q1, we validate the complete PEPA loop in a real indoor office environment requiring sustained autonomous operation across multiple floors. Sys1 provides two inter-floor locomotion primitives. Elevator navigation models the interaction as a four-phase finite state machine: call, wait, enter and select floor, and exit with relocalization; panel approach combines coarse global navigation with precise local alignment, and button presses are executed via visual servoing [28, 24]. Staircase navigation builds on PCT-planner [31] and adds a height-aligned costmap that computes $\Delta z=z(\mathbf{p})-z(\mathbf{r})$ with respect to the nearest waypoint on the global path, enabling the DWA local controller to operate on inclined surfaces. Table I confirms 100% success on both ascent and descent versus 0% for a fixed-height baseline. Code is publicly available. The complete action space comprises 19 discrete actions: idle, eight locomotion primitives, eight emotional expressions (e.g., express_happy), and move_to(location, method).
Fig. 3 shows the robot operating under the Day 3 policy with all three systems active. (a)–(b) demonstrate elevator use as a cross-floor skill executed by Sys1. (c)–(d) show staircase traversal on ascent and descent. (e) shows personality-driven emotional expression, where Sys3’s intrinsic reward shapes the behavior through Sys2 action selection. (f) shows autonomous charging triggered by Sys3’s self-preservation goal, confirming that Sys2 and Sys3 are active alongside Sys1 in real-world operation. Please refer to the submitted multimedia or webpage for more details.
TABLE I: Ablation on local costmap construction for staircase navigation (10 real-world trials: 5 ascents, 5 descents).
@@ IV-A2 Skills and Real-World Deployment
Fixed-height slicing 0/5 0/5 0/10
Height-aligned (ours) 5/5 5/5 10/10
@@ IV-B Personality-Driven Behavior Experiments
We represent personality using natural language descriptions (three sentences define one robot dog), which Sys3 interprets and maps to intrinsic reward components. In PEPA, Sys3 is driven by Qwen3-235B [29] hosted on cloud. The three sentences are the sole personality input to Sys3 and no structured encoding is passed to the model. Full definitions for all five personalities are available at the project website. Drawing from the Big Five personality model, we focus on three behaviorally meaningful dimensions: exploration tendency (adapted from Openness), risk avoidance (adapted from Neuroticism), and goal orientation (adapted from Conscientiousness). To enable controlled comparison, we define five prototype personalities as interpretable anchors in the continuous personality space. The Big Five coordinates serve only as an analytical label to distinguish the prototypes. They are not provided to Sys3. The prototypes span distinct regions of the Big Five trait space (denoted as Openness/Neuroticism/Conscientiousness): Lazy (Low/Med-High/Low); Playful (High/Med/Med); Cautious (Low/High/Med-High); Working (Low/Med/High); and Curious (High/Med/Low-Med). For the full description of 5 personalities, please refer to the project webpage.
Due to the high cost and safety risks of extended physical trials, we evaluate in a sandbox simulator. The simulator uses hand-crafted rules calibrated from real hardware measurements: each action type deducts a fixed battery cost; motor temperature rises with each non-resting action and recovers during rest; and charging replenishes battery at a fixed rate over time. A 0% battery level triggers episode failure, modeling the safety constraint of the physical platform. Each experiment follows a three-day iterative loop.
@@
Sys3 generates an ultimate goal (long-horizon personality anchor) and daily goals that shape intrinsic rewards. For the details of system prompt, please refer to the project webpage. Table II shows representative goals for Day2 and Day3, demonstrating how daily goals evolve through memory-driven reflection. For example, the Playful personality progressively tightens its exploration threshold from 35% to 40% battery after experiencing near-failures. The continual updating of daily goals based on accumulated experience illustrates the agent’s capacity for self-evolution, realizing the open-ended behavioral development central to our framework.
@@
Fig. 5 shows survival statistics across three days. On Day 1, all personalities exhaust batteries and fail. Facing these successive survival failures, Sys3 conducts memory-driven reflection based on episodic experience, reweights daily goal priorities, and adaptively corrects the reward function. This mechanism suppresses extreme personality-biased behaviors and shifts more actions toward self-preservation, such as energy-saving movement and reasonable power management. With such iterative adjustment, Day 3 achieves full survival for all personalities with 72%–100% remaining battery. This result reveals that the optimized personality-conditioned rewards can be effectively co-optimized with self-preservation constraints via continuous memory iteration and strategic adaptation.
IV-C Comparison with Single-LLM Architectures
To answer Q3, we design two baselines that collapse our three-layer architecture into one. The all-cloud baseline uses a powerful cloud-hosted LLM (Qwen3-235B [29]) to handle the entire pipeline, including personality conditioning, daily goal setting, and per-step action decisions, without any computational constraint. The all-edge baseline handles the same full pipeline but is restricted entirely to on-device computation, using a small model (Qwen2.5-1.5B [30]) to eliminate cloud dependency.
TABLE IV: Architecture comparison results. All-cloud matches PEPA in survival time but suffers from high inference latency. All-edge maintains real-time speed but barely improves survival due to the limited reasoning capacity of the small on-device model. PEPA achieves both real-time on-device inference and full 24 h survival.
All-cloud All-edge PEPA
Cloud LLM calls / day 320.6 0 1
Inference latency (s) 25.64 3.54 0.15
Day 3 avg. survival (h) 24 2.85 24
Deployment Cloud On-device Hybrid
Table IV summarizes the results. The all-cloud baseline achieves Day 3 survival comparable to PEPA, confirming that the distilled Sys2 model successfully transfers the decision quality of a large LLM to the edge. However, it requires 320.6 cloud calls per day at 25.64 s each, incompatible with real-time robot operation. The all-edge baseline eliminates cloud dependency but produces no consistent improvement in survival time across days (Day 3 average: 2.85 h), because the 1.5B model lacks the reasoning capacity for effective iterative reflection. PEPA combines the strengths of both: cloud Sys3 is invoked once per day for high-quality reflection, while on-device Sys2 handles real-time action decisions.
IV-D Action Preference Under Identical States
@@ V Conclusion
We present PEPA, a framework that enables long-lived, personality-driven self-evolution in physical agents. Through a three-layer cognitive architecture, we demonstrate that personality can function as an intrinsic driver of behavior, enabling agents to maintain coherent character while adapting to accumulated experiences under real-world constraints. Experimental validation on a quadruped robot platform confirms that distinct personality prototypes exhibit stable, distinguishable behavioral patterns, and that memory-driven reflection progressively improves personality alignment while co-optimizing safety constraints. A primary limitation of the current system is that Sys1 provides a fixed, finite action repertoire, which bounds the behavioral novelty that Sys3’s goal refinement can express in response to unexpected situations. Enabling Sys1 to autonomously acquire new action primitives from experience is a key direction that would allow the full self-evolution loop to operate without pre-specified behavioral constraints.